# GenAI Arena: An Open Evaluation Platform for Generative Models

### v2

@@ Abstract
Generative AI has made remarkable strides to revolutionize fields such as image and video generation. These advancements are driven by innovative algorithms, architecture, and data. However, the rapid proliferation of generative models has highlighted a critical gap: the absence of trustworthy evaluation metrics. Current automatic assessments such as FID, CLIP, FVD, etc often fail to capture the nuanced quality and user satisfaction associated with generative outputs. This paper proposes an open platform GenAI-Arena to evaluate different image and video generative models, where users can actively participate in evaluating these models. By leveraging collective user feedback and votes, GenAI-Arena aims to provide a more democratic and accurate measure of model performance. It covers three arenas for text-to-image generation, text-to-video generation, and image editing respectively. Currently, we cover a total of 27 open-source generative models. GenAI-Arena has been operating for four months, amassing over 6000 votes from the community. We describe our platform, analyze the data, and explain the statistical methods for ranking the models. To further promote the research in building model-based evaluation metrics, we release a cleaned version of our preference data for the three tasks, namely GenAI-Bench. We prompt the existing multi-modal models like Gemini, GPT-4o to mimic human voting. We compute the correlation between model voting with human voting to understand their judging abilities. Our results show existing multimodal models are still lagging in assessing the generated visual content, even the best model GPT-4o only achieves a Pearson correlation of 0.22 in the quality subscore, and behaves like random guessing in others.
@@ 1 Introduction
Automatically assessing the quality of generated visual content is a challenging problem for several reasons: (1) images and videos have many different aspects like visual quality, consistency, alignment, artifacts, etc. Such a multi-faceted nature makes the evaluation intrinsically difficult. (2) the supervised data is relatively scarce on the web. In our work, we release the user voting data as GenAI-Bench to enable further development in this field. Specifically, we calculate the correlation between different image/video auto-raters (i.e. MLLM judges like GPT-4o, Gemini, etc.) with user preference to understand their judging abilities. Our results show that even the best MLLM, GPT-4o achieves at most 0.22 Pearson correlation with human preference.
@@ 5.2 GenAI-Bench Leaderboard
Table 6: GenAI-Bench leaderboard results. It’s designed to benchmark MLLMs’s ability in judging the quality of AI generative contents by comparing with human preferences.
@@ 6 Conclusion
In this paper, we introduced GenAI-Arena, an open platform designed to rank generative models across text-to-image, image editing, and text-to-video tasks based on user preference. unlike other platforms, GenAI-Arena is driven by community voting to ensure transparency and sustainable operation. We employed the side-by-side human voting method to evaluate the models and collected over 6000 votes starting from February 11th, 2024. We compiled an Elo leaderboard with the votings and found that PlayGround V2.5, MagicBrush, and T2V-Turbo are the current state-of-the-art models in the three tasks (until June 4th, 2024). Analysis based on the collected votes shows that while the Elo rating is generally functional, but can biased by the imbalance of the "easy games" and "hard games". Several case studies demonstrate the high quality of our collected votes. What’s more, we also released the human preference voting as GenAI-Bench. We prompt the existing MLLMs to evaluate the generated images and videos on GenAI-Bench and compute the correlations with human voting. The experiment showed that the existing MLLMs achieve very low correlation, even the best model GPT-4o can only achieve $0.22$ Pearson correlation with human voting on quality, same as random guessing on other aspects. In the future, we will continue collecting human votes to update the leaderboard, helping the community to keep track of the research progress. We also plan to develop a more robust MLLM to better approximate human ratings in GenAI-Bench.

### v3

@@ Abstract
Generative AI has made remarkable strides to revolutionize fields such as image and video generation. These advancements are driven by innovative algorithms, architecture, and data. However, the rapid proliferation of generative models has highlighted a critical gap: the absence of trustworthy evaluation metrics. Current automatic assessments such as FID, CLIP, FVD, etc often fail to capture the nuanced quality and user satisfaction associated with generative outputs. This paper proposes an open platform GenAI-Arena to evaluate different image and video generative models, where users can actively participate in evaluating these models. By leveraging collective user feedback and votes, GenAI-Arena aims to provide a more democratic and accurate measure of model performance. It covers three arenas for text-to-image generation, text-to-video generation, and image editing respectively. Currently, we cover a total of 27 open-source generative models. GenAI-Arena has been operating for four months, amassing over 6000 votes from the community. We describe our platform, analyze the data, and explain the statistical methods for ranking the models. To further promote the research in building model-based evaluation metrics, we release a cleaned version of our preference data for the three tasks, namely GenAI-Bench. We prompt the existing multi-modal models like Gemini, GPT-4o to mimic human voting. We compute the accuracy by comparing the model voting with the human voting to understand their juding abilities. Our results show existing multimodal models are still lagging in assessing the generated visual content, even the best model GPT-4o only achieves an average accuracy of $49.19\%$ across the three generative tasks. Open-source MLLMs perform even worse due to the lack of instruction-following and reasoning ability in the complex vision scenarios.
@@ 1 Introduction
Automatically assessing the quality of generated visual content is a challenging problem for several reasons: (1) images and videos have many different aspects like visual quality, consistency, alignment, artifacts, etc. Such a multi-faceted nature makes the evaluation intrinsically difficult. (2) the supervised data is relatively scarce on the web. In our work, we release the user voting data as GenAI-Bench to enable further development in this field. Specifically, we calculate the accuracy between different image/video auto-raters (i.e. MLLM judges like GPT-4o, Gemini, etc.) with user preference to understand their judging abilities. Our results show that even the best MLLM, GPT-4o achieves at most $49.19\%$ accuracy compared with human preference.
@@ 5.2 GenAI-Bench Leaderboard
Table 6: GenAI-Bench leaderboard designed to benchmark MLLMs’s ability in judging the quality of AI generative contents by comparing with human preferences. Numbers are accuracy (%).
@@ 6 Conclusion
In this paper, we introduced GenAI-Arena, an open platform designed to rank generative models across text-to-image, image editing, and text-to-video tasks based on user preference. unlike other platforms, GenAI-Arena is driven by community voting to ensure transparency and sustainable operation. We employed the side-by-side human voting method to evaluate the models and collected over 6000 votes starting from February 11th, 2024. We compiled an Elo leaderboard with the votings and found that PlayGround V2.5, MagicBrush, and T2V-Turbo are the current state-of-the-art models in the three tasks (until June 4th, 2024). Analysis based on the collected votes shows that while the Elo rating is generally functional, but can biased by the imbalance of the "easy games" and "hard games". Several case studies demonstrate the high quality of our collected votes. What’s more, we also released the human preference voting as GenAI-Bench. We prompt the existing MLLMs to evaluate the generated images and videos on GenAI-Bench and compute the accuracy with human voting. The experiment showed that the open-source MLLMs achieve very low performance, even the best model GPT-4o can only achieve $49.19\%$ accuracy. This is mostly because their lack of instruction-following and reasoning ability in complex vision scenarios. In the future, we will continue collecting human votes to update the leaderboard, helping the community to keep track of the research progress. We also plan to develop a more robust MLLM to better approximate human ratings in GenAI-Bench.